OpenStudy (anonymous):

5^-3 ? can someone please explain

OpenStudy (anonymous):

1/125

OpenStudy (anonymous):

1/5^3

OpenStudy (saifoo.khan):

\[\frac{1}{5^3}\]\[\frac{1}{5*5*5}\]\[\frac{1}{125}\]

OpenStudy (anonymous):

the negative sign indicates that the exponent will be a fraction. 1/5^3, which would be 1/125

OpenStudy (anonymous):

1/5^3 1/5∗5∗5 1/125